OpenStudy (anonymous):

the sides of a rectangle is the square root of 14 and the square root of 20.. what is the area?

OpenStudy (anonymous):

Area of a rectangle = length x width. The length could be sqrt(14) and the width could be sqrt(20). All we have to do is multiply the two radicals together. \[A = l*w=\sqrt{14}\sqrt{20}=\sqrt{280}=2\sqrt{70}\]